# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ Introduction
In Section˜7.2, we apply this theorem to get simplest cases of the monopole operators, which are rational versions of certain degenerations of Macdonald difference operators, as the images of particular minor functions on $\mathrm{GL}_{N}$. We express the latter in terms of the shuffle algebra similarly to the geometric and $q$-setting of [FT19] and [Tsy23]. Finally, in Section˜7.1, we show that this isomorphism coincides with the one in [BFN19, Appendix B] under the derived Satake isomorphism of [BF08] between $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ and the equivariant Borel-Moore homology $H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr})$ of the affine Grassmannian of $\mathrm{GL}_{N}$.
We would like to mention that we actively use the formulas (adapted to setting of the paper) from the paper [Mol02] that studied Olshanski centralizer construction from the point of view of Micklesson algebras whose particular case we recall in Section˜4. It would be interesting to find a direct conceptual link between two subjects.
@@
In Section˜1, we introduce the necessary notations concerning $\mathrm{GL}_{N}$ and differential operators $\mathrm{D}(\mathrm{GL}_{N})$ as well quantum minors. In Section˜2, we recall basic properties of shifted Yangians, degenerate affine Hecke algebras, and relate the latter to Shibukawa-Ueno $R$-operators and shuffle algebras. In Section˜3, we recall the definition of Harish-Chandra bimodules and degenerate affine Hecke algebras action on the free ones associated to tensor powers of the vector representation. In Section˜4, we study quantum Hamiltonian reduction with a trivial character, the parabolic restriction. In Section˜5, we study the Kostant-Whittaker reduction and its mirabolic version relating them to degenerate affine Hecke algebras, shuffle algebras, and (the positive part of) shifted Yangians. In Section˜6, we relate two reductions via finite Miura transform and interpret it as a GKLO map. In Section˜7, we prove an isomorphism between truncated shifted Yangians and Toda lattice, compare it with the geometric identification via homology of the affine Grassmannian, and interpret monopole operators in terms certain minor functions. In Appendix˜A, we prove some technical lemmas from the main sections. While in the main body of the paper we deal with the vector representation, in Appendix˜B, we present the corresponding facts for the dual that are needed for the full Toda lattice.
@@
Acknowledgements
I would like to thank Leonid Rybnikov, Joel Kamnitzer, Alexis Leroux-Lapierre, and Théo Pinet for helpful discussions and explanations. This work was supported by Fondation Courtois.
1. Preliminaries
1.1. Notations for $\mathrm{GL}_{N}$
@@
Let $\mathbf{C}^{N}$ be the vector space with a basis $\{v_{1},\ldots,v_{N}\}$. Denote by $\{\phi_{1},\ldots,\phi_{N}\}$ the dual basis of $(\mathbf{C}^{N})^{*}$ such that $\langle v_{i},\phi_{j}\rangle=\delta_{ij}$. Let $\mathrm{Mat}_{N\times N}$ be the affine space of $N\times N$-matrices. We identify its space of functions with the polynomial algebra via
$\mathcal{O}(\mathrm{Mat}_{N\times N})\cong\mathbf{C}[x_{ij}|i,j=1\ldots N],\qquad x_{ij}(A)=\langle A\cdot v_{j},\phi_{i}\rangle,\ A\in\mathrm{Mat}_{N\times N},$
and we combine the generators into a matrix $X\in\mathbf{C}[x_{ij}]\otimes\mathrm{End}(\mathbf{C}^{N})$ via $X_{ij}=x_{ji}$ (observe the transposition).
Let $\mathrm{GL}_{N}\subset\mathrm{Mat}_{N\times N}$ be the group of invertible matrices. Its ring of functions is the localization of $\mathcal{O}(\mathrm{Mat}_{N\times N})$ by the determinant function, in particular, we have an embedding $\mathbf{C}[x_{ij}]\hookrightarrow\mathcal{O}(\mathrm{GL}_{N})$. We call the image the polynomial functions. In addition, there are entries of the matrix inverse $X^{-1}\in\mathcal{O}(\mathrm{GL}_{N})\otimes\mathrm{End}(\mathbf{C}^{N})$ such that
$(X^{-1})_{ij}(A)=\langle A\phi_{j},v_{i}\rangle=\langle A^{-1}v_{i},\phi_{j}\rangle.$
@@
$(g^{R}\cdot f)(A)=f(Ag),\qquad(g^{L}\cdot f)(A)=f(g^{-1}A).$
In particular, for every $\alpha=1\ldots N$, we have equivariant embeddings
(1.1) $\begin{split}{}^{R}\iota_{\alpha}\colon S^{\bullet}\mathbf{C}^{N}&\rightarrow\mathcal{O}(\mathrm{GL}_{N}),\quad v_{i}\mapsto X_{i\alpha},\\ {}^{R}\iota_{\alpha}^{*}\colon S^{\bullet}(\mathbf{C}^{N})^{*}&\rightarrow\mathcal{O}(\mathrm{GL}_{N}),\quad\phi_{i}\mapsto(X^{-1})_{\alpha i}\end{split}$
@@
(1.2) $\begin{split}{}^{L}\iota_{\alpha}\colon S^{\bullet}\mathbf{C}^{N}&\rightarrow\mathcal{O}(\mathrm{GL}_{N}),\quad v_{i}\mapsto(X^{-1})_{i\alpha},\\ {}^{L}\iota_{\alpha}^{*}\colon S^{\bullet}(\mathbf{C}^{N})^{*}&\rightarrow\mathcal{O}(\mathrm{GL}_{N}),\quad\phi_{i}\mapsto X_{\alpha i},\end{split}$
@@
$\mathbb{Z}^{N}=\mathrm{span}_{\mathbb{Z}}(\epsilon_{1},\ldots,\epsilon_{N})\xrightarrow{\sim}\Lambda,\qquad\epsilon_{i}(\mathrm{diag}(t_{1},\ldots,t_{N}))=t_{i}.$
@@
$\Lambda^{+}=\{(\mu_{1},\ldots,\mu_{N})\in\mathbb{Z}^{N}|\mu_{i}\geq\mu_{i+1},i=1\ldots N-1\},$
@@
$\omega_{i}:=(\underbrace{1,\ldots,1}_{i},0,\ldots,0).$
@@
(1.3) $\Delta=\{\epsilon_{i}-\epsilon_{j}|i\neq j\}\subset\Lambda,\qquad\Delta_{+}=\{\epsilon_{i}-\epsilon_{j}|i<j\}\subset\Lambda^{+}$
@@
(1.4) $\rho_{N}=\frac{1}{2}\sum_{\alpha\in\Delta_{+}}\alpha=\sum_{i}\omega_{i}=(N-1,\ldots,1,0)$
@@
$V_{(1,0,\ldots,0)}$ $=\mathbf{C}^{N},$ $\qquad v_{(1,0,\ldots,0)}$ $=v_{1},$
$V_{(0,\ldots,0,-1)}$ $=(\mathbf{C}^{N})^{*},$ $\qquad v_{(0,\ldots,0,-1)}$ $=\phi_{N},$
We identify its Lie algebra $\mathfrak{gl}_{N}$ with $\mathrm{End}(\mathbf{C}^{N})$ via matrix units such that the action on $\mathbf{C}^{N}$ and $(\mathbf{C}^{N})^{*}$ is given by
$\mathfrak{gl}_{N}=\mathrm{span}(E_{ij}|i,j=1\ldots N),\qquad E_{ij}\cdot v_{k}=\delta_{jk}v_{i},\qquad E_{ij}\cdot\phi_{k}=-\delta_{ik}\phi_{j}.$
We combine the generators into a matrix $E\in\mathfrak{gl}_{N}\otimes\mathrm{End}(\mathbf{C}^{N})$:
(1.5) $E=\begin{pmatrix}E_{11}&\ldots&E_{1,N-1}&E_{1N}\\ \vdots&\ddots&\vdots&\vdots\\ E_{N-1,1}&\ldots&E_{N-1,N-1}&E_{N-1,N}\\ E_{N1}&\ldots&E_{N,N-1}&E_{NN}\end{pmatrix}$,
The Lie algebras of $N_{+},N_{-},H_{N}$ are respectively
$\mathfrak{n}_{+}=\mathrm{span}(E_{ij}|i<j),\qquad\mathfrak{n}_{-}=\mathrm{span}(E_{ij}|i>j),\qquad\mathfrak{h}_{N}=\mathrm{span}(E_{ii}).$
@@
(1.6) $\varpi_{N}\colon\mathfrak{gl}_{N}\rightarrow\mathfrak{gl}^{\mathrm{opp}}_{N},\qquad E_{ij}\mapsto E_{ji},$
@@
(1.7) $\tilde{\varpi}_{N}\colon\mathfrak{gl}_{N}\rightarrow\mathfrak{gl}_{N},\qquad E_{ij}\mapsto-E_{ji}.$
Obviously, in terms of the matrix $E$, we have $\varpi(E)=E^{T}$.
1.2. Quantum minors
Denote by
(1.8) $T(u)=u+E\in\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})[u]\otimes\mathrm{End}(\mathbf{C}^{N})$
the T-operator.
Let $(a_{1},\ldots a_{m}),(b_{1},\ldots,b_{m})$ be two sequences of indices. Define a quantum minor as the sum over all permutations of $1,2,\ldots,m$, see [Mol07, (1.54),(1.55)]:
(1.9) $\begin{split}T_{b_{1}\ldots b_{m}}^{a_{1}\ldots a_{m}}(u):=&\sum\limits_{\sigma}\mathrm{sgn}(\sigma)T_{a_{\sigma(1)}b_{1}}(u)T_{a_{\sigma(2)},b_{2}}(u-\hbar)\ldots T_{a_{\sigma(m)}b_{m}}(u-m\hbar+\hbar)=\\ =&\sum\limits_{\sigma}\mathrm{sgn}(\sigma)T_{a_{1}b_{\sigma(1)}}(u-m\hbar+\hbar)\ldots T_{a_{m-1}b_{\sigma(m-1)}}(u-\hbar)T_{a_{m}b_{\sigma(m)}}(u).\end{split}$

### v4

@@ 1 Introduction
In Section 8.2, we apply this theorem to get simplest cases of the monopole operators, which are rational versions of certain degenerations of Macdonald difference operators, as the images of particular minor functions on $\mathrm{GL}_{N}$. We express the latter in terms of the shuffle algebra similarly to the geometric and $q$-setting of [16] and [46]. Finally, in Section 8.1, we show that this isomorphism coincides with the one in [4, Appendix B] under the derived Satake isomorphism of [3] between $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ and the equivariant Borel–Moore homology $H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr})$ of the affine Grassmannian of $\mathrm{GL}_{N}$.
We would like to mention that we actively use the formulas (adapted to setting of the paper) from the paper [37] that studied Olshanski centralizer construction from the point of view of Mickelsson algebras whose particular case we recall in Section 5. It would be interesting to find a direct conceptual link between two subjects.
Future directions
The results of the paper should be more or less directly generalizable to the trigonometric, or $q$-deformed, case: for instance, see [16] for geometric approach, [5] for a certain version of $K$-theoretic Satake equivalence, and [28] and references therein for quantum Harish-Chandra bimodules and the corresponding parabolic restriction. In fact, the only missing ingredient is a $q$-analog of the Kostant–Whittaker reduction which, to the best of our knowledge, was not yet defined; however, see [43] for the $q$-analog of the quantum Toda lattice.
Our original motivation, however, was the elliptic version, where various representation-theoretic constructions should acquire much clearer algebro-geometric meaning, but, unfortunately, much less is known. For instance, to the best of our knowledge, there is no definition of “elliptic” Harish-Chandra bimodules, although there is a strong evidence that such a category should exist, see the discussion in the introduction of [28]. We believe that the results of this paper can give a hint how to approach this problem at least in type A. For example, as the results of Section 6 combined with calculations with Shibukawa–Ueno $R$-operators of Section 3 suggest, the action of degenerate affine Hecke algebras plays a crucial role to study the category of Harish-Chandra bimodules, therefore, it is reasonable to search for their elliptic generalization in terms of representations of elliptic affine Hecke algebras [48]. In general, as Remark 6.15 indicates, it would be interesting to study further this interplay between representation theory and, in a certain sense, “quantization” of various algebro-geometric constructions.
@@
In Section 2, we introduce the necessary notations concerning $\mathrm{GL}_{N}$ and differential operators $\mathrm{D}(\mathrm{GL}_{N})$ as well quantum minors. In Section 3, we recall basic properties of shifted Yangians, degenerate affine Hecke algebras, and relate the latter to Shibukawa–Ueno $R$-operators and shuffle algebras. In Section 4, we recall the definition of Harish-Chandra bimodules and degenerate affine Hecke algebras action on the free ones associated to tensor powers of the vector representation. In Section 5, we study quantum Hamiltonian reduction with a trivial character, the parabolic restriction. In Section 6, we study the Kostant–Whittaker reduction and its mirabolic version relating them to degenerate affine Hecke algebras, shuffle algebras, and (the positive part of) shifted Yangians. In Section 7, we relate two reductions via finite Miura transform and interpret it as a GKLO map. In Section 8, we prove an isomorphism between truncated shifted Yangians and Toda lattice, compare it with the geometric identification via homology of the affine Grassmannian, and interpret monopole operators in terms certain minor functions. In Appendix A, we prove some technical lemmas from the main sections. While in the main body of the paper we deal with the vector representation, in Appendix B, we present the corresponding facts for the dual that are needed for the full Toda lattice.
@@
List of notations
$v_{i}$ natural basis of $\mathbf{C}^{N}$
$\phi_{i}$ dual basis of $\mathbf{C}^{N}$
$X=(x_{ji})$ the transposed matrix of matrix entries functions on $\mathrm{GL}_{N}$
$\mathbf{Z}_{\mathfrak{gl}_{N}}$ center of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$
$A(u)$ generating polynomial of $Z_{\mathfrak{gl}_{N}}$, certain quantum determinant
$\mathfrak{H}^{\kappa}_{k}$ degenerate affine Hecke algebra with quantization parameter $\kappa=\pm\hbar$
$Y_{i}$ polynomial generators of $\mathfrak{H}^{\kappa}_{k}$
$\sigma_{i}$ simple transposition in $\mathbf{C}[S_{k}]\subset\mathfrak{H}_{k}^{\kappa}$
$e_{k}$ symmetrizer in $\mathbf{C}[S_{k}]$
$e_{k}^{-}$ antisymmetrizer in $\mathbf{C}[S_{k}]$
$\sigma^{\kappa}_{i}$ action of $\sigma_{i}$ on spherical module $\mathbf{C}[Y_{1},\dots,Y_{k}]$ of $\mathfrak{H}^{\kappa}_{k}$
$\mathbf{C}[\hbar][Y_{1},\dots,Y_{k}]_{<N}$ polynomials of degree less than $N$ in each variable
$\Omega_{i}$ action of $Y_{i}\in\mathfrak{H}^{\hbar}_{k}$ on Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$
$w_{i}$ shifted coordinates $E_{ii}+N\hbar$-$\hbar$ on $\mathfrak{h}_{N}^{*}\times\mathbb{A}^{1}$
$\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$ localized ring of functions on $\mathfrak{h}_{N}^{*}\times\mathbb{A}^{1}$
$P$ extremal projector of $\mathfrak{gl}_{N}$
$^{N}\!\tilde{\mathcal{A}}_{\hbar}$ algebra of localized difference operators $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\bigl[\mathbf{u}_{1}^{\pm},\dots,\mathbf{u}_{N}^{\pm}\bigr]$
$^{N}\!\tilde{\mathcal{A}}_{\hbar}^{-}$ negative part of difference operators $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\bigl[\mathbf{u}_{1}^{-1},\dots,\mathbf{u}_{N}^{-1}\bigr]$
$\psi$ non-zero number in $\mathbf{C}$
$\psi$ character of $\mathfrak{n}_{+}$ such that $\psi(E_{ij})=\delta_{i+1,j}\psi$
$\mathfrak{m}_{N}$ mirabolic subalgebra
$()^{R},()^{L}$ anything related to right (resp. left) action of $\mathrm{GL}_{N}$ on itself
$^{R}\iota_{\alpha}$ and analogs map $\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}^{R}_{N}\bigr)\ltimes S^{\bullet}\mathbf{C}^{N}\rightarrow\mathrm{D}_{\hbar}(\mathrm{GL}_{N})$ or their reductions
$\hat{T}(u)$ quantum comatrix of $T(u)=u+E$
$v_{\lambda},v_{\lambda}^{*}$ highest-weight vector in irreducible representation $V_{\lambda}$ (resp. its dual $V_{\lambda}^{*}$)
$\mathrm{Gr}$ affine Grassmannian $\mathrm{GL}_{N}((t))/\mathrm{GL}_{N}[[t]]$
$\mathrm{Gr}^{\lambda}$ $\mathrm{GL}_{N}[[t]]$-orbit of $t^{\lambda}\in\mathrm{Gr}$
$\Omega_{i}^{*}$ action of $Y_{i}\in\mathfrak{H}_{k}^{-\hbar}$ on Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\bigl(\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)^{\otimes k}$
2 Preliminaries
2.1 Notations for $\boldsymbol{\mathrm{GL}_{N}}$
@@
Let $\mathbf{C}^{N}$ be the vector space with a basis $\{v_{1},\dots,v_{N}\}$. Denote by $\{\phi_{1},\dots,\phi_{N}\}$ the dual basis of $\bigl(\mathbf{C}^{N}\bigr)^{*}$ such that $\langle v_{i},\phi_{j}\rangle=\delta_{ij}$. Let $\mathrm{Mat}_{N\times N}$ be the affine space of $N\times N$ matrices. We identify its space of functions with the polynomial algebra via
$\mathcal{O}(\mathrm{Mat}_{N\times N})\cong\mathbf{C}[x_{ij}\mid i,j=1,\dots,N],\qquad x_{ij}(A)=\langle A\cdot v_{j},\phi_{i}\rangle,\quad A\in\mathrm{Mat}_{N\times N},$
and we combine the generators into a matrix
$X=(x_{ji})_{i=1,\dots,N}^{j=1,\dots,N}\in\mathbf{C}[x_{ij}]\otimes\mathrm{End}\bigl(\mathbf{C}^{N}\bigr)$
(observe the transposition).
Let $\mathrm{GL}_{N}\subset\mathrm{Mat}_{N\times N}$ be the group of invertible matrices. Its ring of functions is the localization of $\mathcal{O}(\mathrm{Mat}_{N\times N})$ by the determinant function, in particular, we have an embedding $\mathbf{C}[x_{ij}]\hookrightarrow\mathcal{O}(\mathrm{GL}_{N})$. We call the image the polynomial functions. In addition, there are entries of the matrix inverse $X^{-1}\in\mathcal{O}(\mathrm{GL}_{N})\otimes\mathrm{End}\bigl(\mathbf{C}^{N}\bigr)$ such that
$\bigl(X^{-1}\bigr)_{ij}(A)=\langle v_{i},A\phi_{j}\rangle=\bigl\langle A^{-1}v_{i},\phi_{j}\bigr\rangle.$
@@
$\bigl(g^{R}\cdot f\bigr)(A)=f(Ag),\qquad\bigl(g^{L}\cdot f\bigr)(A)=f\bigl(g^{-1}A\bigr).$
In particular, for every $\alpha=1,\dots,N$, we have equivariant embeddings
$\begin{split}&{}^{R}\iota_{\alpha}\colon\ S^{\bullet}\mathbf{C}^{N}\rightarrow\mathcal{O}(\mathrm{GL}_{N}),\qquad v_{i}\mapsto X_{i\alpha},\\ &{}^{R}\iota_{\alpha}^{*}\colon\ S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}\rightarrow\mathcal{O}(\mathrm{GL}_{N}),\qquad\phi_{i}\mapsto\bigl(X^{-1}\bigr)_{\alpha i}\end{split}$ (2.1)
@@
$\begin{split}&{}^{L}\iota_{\alpha}\colon\ S^{\bullet}\mathbf{C}^{N}\rightarrow\mathcal{O}(\mathrm{GL}_{N}),\qquad v_{i}\mapsto\bigl(X^{-1}\bigr)_{i\alpha},\\ &{}^{L}\iota_{\alpha}^{*}\colon\ S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}\rightarrow\mathcal{O}(\mathrm{GL}_{N}),\qquad\phi_{i}\mapsto X_{\alpha i},\end{split}$ (2.2)
@@
$\mathbb{Z}^{N}=\mathrm{span}_{\mathbb{Z}}(\epsilon_{1},\dots,\epsilon_{N})\xrightarrow{\sim}\Lambda,\qquad\epsilon_{i}(\mathrm{diag}(t_{1},\dots,t_{N}))=t_{i}.$
@@
$\Lambda^{+}=\bigl\{(\mu_{1},\dots,\mu_{N})\in\mathbb{Z}^{N}\mid\mu_{i}\geq\mu_{i+1},\,i=1,\dots N-1,\bigr\},$
@@
$\omega_{i}:=(\underbrace{1,\dots,1}_{i},0,\dots,0).$
@@
$\Delta=\{\epsilon_{i}-\epsilon_{j}\mid i\neq j\}\subset\Lambda,\qquad\Delta_{+}=\{\epsilon_{i}-\epsilon_{j}\mid i<j\}\subset\Lambda^{+}$
@@
$\rho_{N}=\frac{1}{2}\sum_{\alpha\in\Delta_{+}}\alpha=\sum_{i}\omega_{i}=(N-1,\dots,1,0)$ (2.3)
@@
$V_{(1,0,\dots,0)}=\mathbf{C}^{N},$ $\qquad v_{(1,0,\dots,0)}=v_{1},$
$V_{(0,\dots,0,-1)}=\bigl(\mathbf{C}^{N}\bigr)^{*},$ $\qquad v_{(0,\dots,0,-1)}=\phi_{N},$
We identify its Lie algebra $\mathfrak{gl}_{N}$ with $\mathrm{End}\bigl(\mathbf{C}^{N}\bigr)$ via matrix units such that the action on $\mathbf{C}^{N}$ and $\bigl(\mathbf{C}^{N}\bigr)^{*}$ is given by
$\mathfrak{gl}_{N}=\mathrm{span}(E_{ij}\mid i,j=1,\dots,N),\qquad E_{ij}\cdot v_{k}=\delta_{jk}v_{i},\qquad E_{ij}\cdot\phi_{k}=-\delta_{ik}\phi_{j}.$
We denote by the same letters the corresponding generators of the (asymptotic) universal enveloping algebra $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$ in the sense of Definition 4.1; in particular,
$E_{ij}E_{kl}-E_{kl}E_{ij}=\hbar(\delta_{jk}E_{il}-\delta_{li}E_{kj}).$
We combine the generators into a matrix $E\in\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathrm{End}\bigl(\mathbf{C}^{N}\bigr)$:
$E=\begin{pmatrix}E_{11}&\dots&E_{1,N-1}&E_{1N}\\ \vdots&\ddots&\vdots&\vdots\\ E_{N-1,1}&\dots&E_{N-1,N-1}&E_{N-1,N}\\ E_{N1}&\dots&E_{N,N-1}&E_{NN}\end{pmatrix}$, (2.4)
The Lie algebras of $_{+}, NN_{-}$, $H_{N}$ are respectively
$\mathfrak{n}_{+}=\mathrm{span}(E_{ij}\mid i<j),\qquad\mathfrak{n}_{-}=\mathrm{span}(E_{ij}\mid i>j),\qquad\mathfrak{h}_{N}=\mathrm{span}(E_{ii}).$
@@
$\varpi_{N}\colon\ \mathfrak{gl}_{N}\rightarrow\mathfrak{gl}^{\mathrm{opp}}_{N},\qquad E_{ij}\mapsto E_{ji},$ (2.5)
@@
$\tilde{\varpi}_{N}\colon\ \mathfrak{gl}_{N}\rightarrow\mathfrak{gl}_{N},\qquad E_{ij}\mapsto-E_{ji}.$ (2.6)
Obviously, in terms of the matrix $E$, we have $\varpi(E)=E^{\mathsf{T}}$.
2.2 Quantum minors
Denote by $T(u)=u+E\in\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})[u]\otimes\mathrm{End}\bigl(\mathbf{C}^{N}\bigr)$ the T-operator.
Let $(a_{1},\dots,a_{m})$, $(b_{1},\dots,b_{m})$ be two sequences of indices. Define a quantum minor as the sum over all permutations of $1,2,\dots,m$, see [38, formulas (1.54) and (1.55)]:
$T_{b_{1}\dots b_{m}}^{a_{1}\dots a_{m}}(u):=$ $\sum\limits_{\sigma}\mathrm{sgn}(\sigma)T_{a_{\sigma(1)}b_{1}}(u)T_{a_{\sigma(2)},b_{2}}(u-\hbar)\cdots T_{a_{\sigma(m)}b_{m}}(u-m\hbar+\hbar)$
$=$ $\sum\limits_{\sigma}\mathrm{sgn}(\sigma)T_{a_{1}b_{\sigma(1)}}(u-m\hbar+\hbar)\cdots T_{a_{m-1}b_{\sigma(m-1)}}(u-\hbar)T_{a_{m}b_{\sigma(m)}}(u).$ (2.7)